### v1

# Freestyle Sketch-in-the-Loop Image Segmentation

Abstract
In this paper, we expand the domain of sketch research into the field of image segmentation, aiming to establish freehand sketches as a query modality for subjective image segmentation. Our innovative approach introduces a “sketch-in-the-loop” image segmentation framework, enabling the segmentation of visual concepts partially, completely, or in groupings – a truly “freestyle” approach – without the need for a purpose-made dataset (i.e., mask-free). This framework capitalises on the synergy between sketch-based image retrieval (SBIR) models and large-scale pre-trained models (CLIP or DINOv2). The former provides an effective training signal, while fine-tuned versions of the latter execute the subjective segmentation. Additionally, our purpose-made augmentation strategy enhances the versatility of our sketch-guided mask generation, allowing segmentation at multiple granularity levels. Extensive evaluations across diverse benchmark datasets underscore the superior performance of our method in comparison to existing approaches across various evaluation scenarios.
1 Introduction
This paper delves into the realm of image segmentation, but with a unique twist – the incorporation of freehand sketches. Our research vision is to explore the impact of fine-grained subjectivity inherently captured by a sketch [30, 19], on image segmentation. At its core, our aim is to devise a subjective image segmentation framework that empowers users to freely segment any object, be it a partial segment, a complete entity, or a group of interconnected elements – essentially, enabling image segmentation in a “freestyle” manner.
Image segmentation has witnessed significant evolution [14, 39] over time, and we are committed to keeping pace with the current capabilities of image segmentation systems – we also want to operate in an open-vocabulary [57] manner and harness the potential of the latest foundation models. We too fully acknowledge prior segmentation methods that inject subjectivity through textual inputs [85]. The central argument of this paper however revolves around the premise that sketches offer a more precise form of subjectivity – a point convincingly illustrated in LABEL:fig:teaser (bottom).
To elucidate why sketches serve as a superior modality for capturing subjectivity, we direct readers to the extensive sketch literature [10, 19, 86, 25, 74]. In summary, the insights drawn from this large body of work are two-fold: (i) sketches excel in their ability to convey intricate visual details that may be cumbersome or ambiguous when expressed through text [74]. (ii) sketches exhibit a smaller semantic gap to photos, as both share the same visual modality, albeit in an abstract form [30]. This inherent relationship is notably featured in the extensively studied domain of SBIR [20] and its fine-grained variations [7] – a challenge that we also employ as a guiding critic in our training.
The specifics of our vision present significant challenges. First and foremost, we need to operate without any sketch-mask pairing data. The collection of such a dataset is simply unfeasible, especially considering the current status quo within the sketch community still concerns gathering enough sketch-photo pairs [58, 73], let alone pixel-perfect masks. This challenge is further exacerbated by our open-vocabulary setting, which essentially renders prior arts inapplicable [31]. Secondly, akin to most prior work in the sketch literature [86, 72, 6, 41], we face a substantial gap between the sketch and photo domains.
We commence by fixating on the ultimate goal: given a sketch, compute corresponding segmentation masks for all applicable gallery photos. Surprisingly, it turns out that all we need to compute this correspondence is the strategic utilisation of (i) a frozen Fine-Grained Sketch-Based Image Retrieval (FG-SBIR) model and (ii) a pre-trained foundation model (e.g., CLIP [68], DINOv2 [60]). The former provides a solid foundation for fine-grained sketch-photo correspondence and also serves as a training signal to fine-tune the latter. Meanwhile, the fine-tuned version of the latter predicts the actual mask for the photo using its corresponding sketch, thereby eliminating the need for manually annotated masks (i.e., mask-free). It follows that these two components are intricately linked by this very hypothesis: a sketch and its matching background-masked image pair (and foreground highlighted) will sit closely within the fixed FG-SBIR joint embedding [72]. This, we hope, is intuitive – the more highlighted (using the computed mask) the foreground object is, the easier it is for its matching sketch to locate. It is through this iterative training process that the segmentation model continually hones its performance.
In generating sketch-guided masks, we use pre-trained DINOv2 [60] for the fine-grained setup (CLIP [68] for category-level). It extracts a spatial feature map from a photo and a global feature vector from a sketch. A cosine similarity computation, followed by bilinear upscaling and differentiable thresholding, yields a pseudo mask used for generating the background-masked image corresponding to the given sketch. To prevent all-white mask predictions and overfitting, a mask-regularisation loss is introduced. Additionally, for part-level segmentation, we employ sketch-partitioning, randomly fragmenting binary sketches into orthogonal splits, enabling model-prediction from a specific split, for more precise part-level segmentation.
In summary, our contributions are (i) we address the challenge of subjective segmentation by introducing a novel approach that enables mask-free image segmentation through the interpretive power of sketches. (ii) we support multi-granularity segmentation, providing users with the flexibility to segment visual concepts at various levels, encompassing entire entities, specific parts, or interconnected groups. (iii) we establish an effective synergy between sketch-based image retrieval models and large-scale pre-trained models. (iv) we propose a purpose-made augmentation strategy that enhances the framework’s versatility, enabling segmentation at multiple granularity levels.
2 Related Works
Sketch for Vision Tasks.Freehand sketches has long established its significance for several visual understanding tasks [8, 86, 72, 19, 41, 10]. Following its success in $2D$ and $3D$ image retrieval [86, 72, 24, 54, 53], sketch as a vision-modality has been used in image generation [41], saliency detection [10], inpainting [87], shape-modelling [56], augmented reality [54], medical image analysis [40], image editing [90, 47], object detection [19], representation learning [22], class-incremental learning [8], etc. To our best knowledge, Sketch-a-Segmenter [31] is the only existing work dealing with sketch-guided segmentation. This method [31] however requires fully-annotated pixel-level masks for training [31], unlike ours. We thus here aim to forward sketch research in another interesting direction by utilising its potential as a weak label for segmentation.
Segmentation in Vision.  Image segmentation [70, 14, 15, 49] can be broadly classified into three categories based on task granularity(i) semantic segmentation [49, 81, 93] – aims to classify and label each pixel into a set of pre-defined classes, (ii) instance segmentation [77, 29] – involves identifying different instances of each object, (iii) panoptic segmentation [84, 35] – combines semantic and instance segmentation to simultaneously assign class labels and identify object instances. Rising popularity of vision-language [68] and generative [69, 36] models further inspired open-vocab segmentation frameworks like GroupViT [83], SegCLIP [52], Segmentation-in-Style [61], ODISE [84], etc. The recent Segment Anything Model (SAM) [39] introduces a zero-shot framework capable of handling diverse segmentation tasks in different image distributions [39]. Its success promoted multiple follow-up works like HQ-SAM [37], PerSAM [91], etc. Our aim here is to introduce the subjectivity of freehand sketches for image segmentation. Moreover, unlike prior methods, our approach segments all similar instances of the query sketch from a large gallery of photos instead of segmenting a single image at hand.
Weakly/Un-Supervised Segmentation.  Weakly supervised frameworks typically use image-level labels [4, 45, 65, 66, 67], scribbles [46, 79], bounding-boxes [38, 63, 21], cue points [5], eye-tracking data [62], web-tags [2], saliency masks [59], etc. as a source for weak-supervision. Such methods can be broadly classified as – (i) one stage and (ii) two stage. While one-stage methods [4, 13, 63, 66] directly use image-level labels for training, they tend to perform lower than two-stage ones, which first generate pseudo segmentation labels via attention maps [94, 76] and then use them for training. In contrast, unsupervised image segmentation can be broadly divided into two categories– (i) generative methods utilise unlabelled images to train specialised image generators [3, 16] or use pre-trained generators to directly generate segmentation masks [55, 80, 1]; (ii) discriminative methods are mainly based on clustering and contrastive learning [34, 82, 78, 92].
3 Revisiting Sketch-based Image Retrieval
Here we revisit the baseline SBIR paradigms in favour of the broader vision community. Baseline category-level [20] and fine-grained [72] SBIR models described here will be used later (Sec. 4.3) to guide the training of our mask-free sketch-based category-level and fine-grained image segmentation models respectively.
Baseline Category-level SBIR.  Provided a query sketch $\mathcal{S}\in\mathbb{R}^{H\times W\times 3}$ of any class, category-level SBIR aims to retrieve a photo $p_{i}^{j}$ from the same class, out of a gallery $\mathcal{G}=\{p_{i}^{j}\}_{i=1}^{N_{j}}{|}_{j=1}^{N_{c}}$ containing images from a total $N_{c}$ classes with $N_{j}$ images per class [72]. Particularly, a backbone feature extractor (separate weights for sketch and photo branches) network is trained to generate $d$-dimensional feature vector $f_{i}=\mathcal{F}(\mathcal{I}):\mathbb{R}^{H\times W\times 3}\rightarrow\mathbb{R}^{d}$ via triplet loss [88] that aims to minimise the distance $\delta(a,b)={||a-b||}_{2}$ between an anchor sketch ($\mathcal{S}$) feature $f_{s}$ and a positive photo ($\mathcal{P}$) feature $f_{p}$ from the same category as $\mathcal{S}$, while increasing that from a negative photo ($\mathcal{N}$) feature $f_{n}$ of a different category. With margin $\mu_{\mathrm{cat}}>0$, triplet loss for category-level SBIR could be formulated as: $\mathcal{L}_{\mathrm{triplet}}^{\mathrm{cat}}=\mathtt{max}\{0,\mu_{\mathrm{cat}}+\delta(f_{s},f_{p})-\delta(f_{s},f_{n})\}$
Baseline Fine-grained (FG) SBIR.  Unlike category-level, cross-category FG-SBIR setup intends to learn a single model capable of fine-grained instance-level matching from multiple ($N_{c}$) classes. Akin to category-level [72], cross-category FG-SBIR framework learns a backbone feature extractor $\mathcal{F}(\cdot)$ (weights shared between sketch and photo branches), via triplet loss $\mathcal{L}^{\mathrm{fine}}_{\mathrm{triplet}}$ (similar to $\mathcal{L}^{\mathrm{cat}}_{\mathrm{triplet}}$) but with hard triplets, where the negative sample is a different instance ($\mathcal{P}^{j}_{k};k\neq i$) of the same class as the anchor sketch ($\mathcal{S}_{i}^{j}$) and its matching photo ($\mathcal{P}_{i}^{j}$) [72]. Furthermore, for learning inter-class discrimination, it uses an $N_{c}$-class classification head with cross-entropy loss ($\mathcal{L}_{\mathrm{class}}$) on the sketch-photo joint embedding space. Thus, the total loss [9] for baseline FG-SBIR is: $\mathcal{L}^{\mathrm{fine}}_{\mathrm{triplet}}+\mathcal{L}_{\mathrm{class}}$.
Figure 1: Examples of fine-grained (left) and category-level (right) image pairs from Sketchy [73] dataset.
Datasets.   Category-level datasets like Sketchy-Extended [73, 48], TU-Berlin [26], Quick!, Draw [27], etc. hold category-level association, where every sketch of a category is associated with all photos of the same category as $\{s_{i}^{j},\{p_{k}^{j}\}_{k=1}^{N_{j}^{P}}\}_{i=1}^{N_{j}^{S}}|_{j=1}^{N_{c}}$ from $N_{c}$ classes with every $j^{th}$ class having $N^{S}_{j}$ sketches and $N^{P}_{j}$ photos. Whereas, FG-SBIR datasets (e.g., Sketchy [73]) contain instance-level sketch-photo pairs as $\{s_{i}^{j},p_{i}^{j}\}_{i=1}^{N_{j}}{|}_{j=1}^{N_{c}}$ from $N_{c}$ classes with every $j^{th}$ class having $N_{j}$ sketch-photo pairs with fine-grained association. Here we use category-level and fine-grained SBIR models [71] trained with Sketchy-Extended [73] and Sketchy [73] datasets respectively.
4 Proposed Methodology
Overview.  Given an image $\mathcal{I}\in\mathbb{R}^{h\times w\times 3}$, state-of-the-art supervised image segmentation models [49, 14, 15] aim to generate a segmentation mask $\mathcal{M}\in\mathbb{R}^{h\times w\times(c+1)}$ via pixel-level classification across a pre-defined set of $c$ classes and a background ($+1$). On the contrary, given a query sketch $\mathcal{S}$, our method generates a binary segmentation mask $\mathcal{M}\in\mathbb{R}^{h\times w\times 1}$, depicting the pixel-locations, wherever the queried-concept appears in any candidate image $\mathcal{I}\in\mathbb{R}^{h\times w\times 3}$, with pixels belonging to that concept marked as $1$ and all the rest as 00. Learning segmentation frameworks via typical supervised training setup however, is impractical given the laborious process of collecting human-annotated ground truth segmentation masks. We therefore aim to utilise the existing sketch-photo dataset [73] to devise a weakly-supervised mask-free image segmenter. Although existing weak supervision signals (query) like tags [2], caption [57], cue points [5], etc. might suffice for category-level vision tasks, they lack in fine-grained cues unlike sketches [10, 19]. Inspired by sketch’s success in multiple fine-grained vision tasks (e.g., retrieval [72], detection [19], generation [41], etc.), we aim to establish sketch as a worthy query-modality for mask-free weakly-supervised image segmentation, both at category-level and fine-grained setup. Furthermore, we also extend our method to part-level segmentation, using the same dataset.
4.1 Problem Definition: Segmentation Granularity
Category Level Segmentation.  In this setup, we aim to predict a binary segmentation mask, denoted as $\mathcal{M}\in\mathbb{R}^{h\times w\times 1}$, based on a query sketch $\mathcal{S}_{c}$ for a specific category, denoted as $c$. $\mathcal{M}$ will distinguish the foreground object (of category $c$) by marking the pixels inside the boundary as $1$, while the background pixels are set to 00. Essentially, when you draw a “camel” this system will generate segmentation masks for any “camel” photos in the entire test gallery.
Fine-grained Segmentation.   Given a query sketch $\mathcal{S}_{f}$ depicting a specific object instance of category $c$, we generate similar segmentation masks $\mathcal{M}\in\mathbb{R}^{h\times w\times 1}$ for any photo in the gallery of category $c$ where the object holds the same shape/pose as that of $\mathcal{S}_{f}$. For instance, drawing a “camel sitting down” should generate masks only for those photos where the camel(s) is(are) “sitting down” in the same shape/pose, not any photo containing “camel” (Fig. 2 (left)).
Part-level Segmentation.  Unlike existing weak supervision signals (e.g., tags [2], cue points [5], etc.), sketch has the ability to model part-level query [32]. Accordingly, given a query sketch $\mathcal{S}_{p}$ depicting a specific part of an object from category $c$, the aim is to generate similar segmentation masks $\mathcal{M}\in\mathbb{R}^{h\times w\times 1}$ depicting only the queried part for any photo in $\mathcal{G}$ of category $c$. For instance, drawing “a camel’s head” should generate segmentation masks showing only the “head” for any “camel” photos in $\mathcal{G}$ (Fig. 2 (right)). Here the part-level query $\mathcal{S}_{p}$ is a region-wise spatial-subset ($\mathcal{S}_{p}\subset\mathcal{S}$) of the full query $\mathcal{S}$.
Figure 2: Fine-grained (left) and part-level (right) segmentation.
4.2 Baseline Weakly Supervised Segmentation
In our setup, we consider the available sketches ($\mathcal{S}$) and photos ($\mathcal{P}$) from $\mathcal{D}_{\mathrm{cat}}$= $\{\mathcal{S}_{i},\mathcal{P}_{i}\}_{i=1}^{M}$ and $\mathcal{D}_{\mathrm{fine}}$=$\{\mathcal{S}_{j},\mathcal{P}_{j}\}_{j=1}^{N}$ with category-level [48] and fine-grained [73] association between them respectively (cf. Sec. 3). We design our baseline sketch-based image segmenter taking inspiration from text-based weakly-supervised segmentation framework [57] that employs patch-wise InfoNCE loss between CLIP’s global textual feature and patch-wise visual features [57] for better text-photo region-wise alignment (instead of using global feature for both as in CLIP training [68]) required for dense pixel-level segmentation task [57]. We however exclude the textual encoder and employ CLIP’s visual encoder to extract patch-wise photo features and global query sketch feature to apply the same loss.
In particular, we consider a pre-trained CLIP [68] ViT-L/14 backbone visual feature extractor $\mathbf{V}(\cdot)$ that splits the input $\mathcal{I}$ into $T$ patches and extracts patch-wise features $\{\mathcal{I}_{p}^{i}\}$ and global $\mathtt{[CLS]}$-token feature $\mathcal{I}_{\mathtt{CLS}}$ as $\mathbf{V}(\mathcal{I})$ = $[\mathcal{I}_{p}^{1},\mathcal{I}_{p}^{2},\cdots,\mathcal{I}_{p}^{T},\mathcal{I}_{\mathtt{CLS}}]\in\mathbb{R}^{(T+1)\times d}$. For a sketch-query $\mathcal{S}$, we use its global feature representation $\mathbf{s}=\mathbf{V}_{g}(\mathcal{S})\in\mathbb{R}^{1\times d}$ using the $\mathtt{[CLS]}$-token feature, while for a candidate photo $\mathcal{P}$, we utilise its patch-wise feature representation $\mathbf{p}=\mathbf{V}_{p}(\mathcal{P})\in\mathbb{R}^{T\times d}$. Although $\mathbf{V}$ is shared between sketch and photo, $\mathbf{V}_{g}$ and $\mathbf{V}_{p}$ denotes global and patch-wise feature extraction respectively. We calculate cosine similarity between $\mathbf{s}$ and $\mathbf{p}$ to obtain a patch-wise correlation vector $\xi\in\mathbb{R}^{T\times 1}$ with $\mathtt{SoftMax}$ normalisation across the patch dimension. Finally, taking the weighted sum across the patches, we obtain the sketch-query aware photo feature as ${\mathbf{\hat{p}}}=\sum_{i=1}^{T}(\xi_{i}\times\mathbf{p}_{i})$. We perform $l_{2}$ normalisation on $\mathbf{s}$ and $\mathbf{\hat{p}}$ followed by calculating InfoNCE between them as:
$\mathcal{L}_{\mathrm{InfoNCE}}=\frac{1}{2k}\sum_{i=1}^{k}\left(\frac{\mathtt{exp}({\mathbf{\hat{p}}_{i}\cdot\mathbf{s}_{i}})}{\sum_{j=1}^{k}\mathtt{exp}({\mathbf{\hat{p}}_{i}\cdot\mathbf{s}_{j}})}+\frac{\mathtt{exp}({\mathbf{\hat{p}}_{i}\cdot\mathbf{s}_{i}})}{\sum_{j=1}^{k}\mathtt{exp}({\mathbf{\hat{p}}_{j}\cdot\mathbf{s}_{i}})}\right)\vspace{-0.1cm}$ (1)
Figure 3: We generate sketch-guided correlation map $\mathcal{C}$ by multiplying the global sketch feature $\mathbf{s}$ with reshaped patch-wise photo feature $\mathbf{p}_{s}$. Bilinear upscaling, followed by differentiable Sigmoid thresholding yields patch-wise correlation mask $\mathcal{M}$, which when multiplied with $\mathcal{P}$ gives masked candidate photo $\mathcal{P}_{\mathrm{mask}}$ (highlighted in grey). Frozen SBIR backbone $\mathcal{F}$ enforces the input sketches to generate masks that segment only the foreground object (queried via $\mathcal{S}$) via $\mathcal{L}_{\mathrm{SBIR}}$. $\mathcal{L}_{\mathrm{infoNCE}}$ between $\mathbf{s}$ and $\mathbf{p}_{s}$ enhances sketch-photo alignment. At every batch, $\mathcal{L}_{\mathrm{unpaired}}$ ensures an all-zero mask for negative samples ($\mathcal{N}$). To avoid overfitting to a trivial all-one mask, we additionally regularise $\mathcal{M}$ via $\mathcal{L}_{\mathrm{reg}}$.
Here, $k$ denotes batchsize. For a photo of size $\mathbb{R}^{224\times 224\times 3}$, the number of patches $T$=$256$ for ViT-L/14 backbone. During testing, we calculate cosine similarity between global sketch query feature $\mathbf{s}$in $\mathbb{R}^{1\times d}$ and patch-wise candidate image features $\mathbf{p}$in $\mathbb{R}^{256\times d}$ to obtain the patch-wise correlation vector $\xi$in $\mathbb{R}^{256\times 1}$ that shows the degree to which each image patch correlates to the query-concept. We apply element-wise $\mathtt{Sigmoid}$ on $\xi$ and reshape its spatial tensor equivalent of size $\mathbb{R}^{16\times 16\times 1}$ followed by bilinearly upscaling it to the original photo size, to get the similarity map of size $\mathbb{R}^{h\times w\times 1}$. Empirically thresholding it with a fixed value ($0.5$) we get the predicted binary segmentation mask.
However, this baseline suffers a few drawbacks – (i) while CLIP [68] works reasonably well on category-level tasks [72], it has limited fine-grained potential [72]. Consequently, for the fine-grained setup, we use a pre-trained DINOv2 [60] ViT-L/14 backbone which showcases better fine-grained cross-modal correspondence [60], and (ii) CLIP’s contrastive learning happens in the embedding-space [68], while segmentation is a pixel-level task. Furthermore, alignment in the latent embedding-space does not guarantee pixel-level semantic alignment. While this embedding-space alignment works well in case of semantic-driven text-based segmentation [57], sketches being more shape-driven [72, 6, 18, 58] perform sub-optimally in this case.
4.3 Sketch-guided Mask Generation
Given a sketch and a candidate image, we extract the global sketch-query feature $\mathbf{s}$=$\mathbf{V}_{g}(\mathcal{S})\in\mathbb{R}^{1\times d}$ and patch-wise candidate image features $\mathbf{p}$=$\mathbf{V}_{p}(\mathcal{P})\in\mathbb{R}^{T\times d}$. We reshape (for brevity) the patch-wise features $\mathbf{p}$ in the equivalent image-space spatial order to get the $\mathbf{p}_{s}$=$\mathtt{Reshape}(\mathbf{p})\in\mathbb{R}^{h^{\prime}\times w^{\prime}\times d}$ spatial tensor where $T$=$h^{\prime}\times w^{\prime}$. We get the sketch-guided correlation map $\mathcal{C}$ by computing cosine similarity between $\mathbf{s}$ and $\mathbf{p}_{s}$ using the sketch as a probe-vector as:
$\small\mathcal{C}(u,v)=\frac{\langle\mathbf{s},\mathbf{p}_{s}(u,v)\rangle}{||\mathbf{s}||~{}||\mathbf{p}_{s}(u,v)||}~{}~{}~{}~{}u=[1,..,h^{\prime}],v=[1,..,w^{\prime}]\vspace{-0.1cm}$ (2)
Now the research question remains as to how can we leverage this sketch-guided correlation map $\mathcal{C}$ in the learning process while backpropagating the training signal via pixel space. One solution could be to bilinearly upscale $\mathcal{C}\in\mathbb{R}^{h^{\prime}\times w^{\prime}\times 1}$ to the original photo resolution ($\mathbb{R}^{h\times w\times 1}$), and threshold it with a fixed value and multiply with the candidate photo to get a masked photo, where ideally the foreground objects (w.r.t. the sketch concept) would remain unmasked and the background regions would be suppressed. Nonetheless, hard thresholding with a fixed value invokes non-differentiability due to the involvement of a Heaviside function [12]. To alleviate this issue, we approximate the Heaviside thresholding [12] operation with a temperature ($\tau$) controlled Sigmoid function [64, 12] as:
$\mathcal{M}={1}\big{/}{(1+\mathtt{exp}((\mathcal{C}-0.5)/\tau))}\vspace{-0.15cm}$ (3)
where $\tau$ is the temperature parameter defining the sharpness of the Sigmoid, and $0.5$ is the thresholding value. With the thresholded patch-wise correlation mask ${\mathcal{M}}$, we generate the masked candidate photo via spatial-wise multiplication as $\mathcal{P}_{\mathrm{mask}}$=$\mathcal{P}\times{\mathcal{M}}$. During training, we aim to jointly optimise the visual encoder such that the mask generates true responses, only in those spatial regions of the photo that contain the object drawn in the input sketch. Now, to quantify the quality of the masked photo in terms of background-foreground separation, we use a pre-trained SBIR [71] model-based loss described in the next section.
4.4 Training Objectives
Pre-trained SBIR Loss.  We hypothesise that the masked photo $\mathcal{P}_{\mathrm{mask}}$ and the query sketch $\mathcal{S}$ will lie very close in the pre-trained SBIR model’s ($\mathcal{F}$) embedding space when the generated mask $\mathcal{M}$ would only segment the foreground object related to the sketch query. However, when the generated mask falsely suppresses the foreground object, $\mathcal{P}_{\mathrm{mask}}$ would likely diverge from $\mathcal{S}$, in the same space. Therefore, $\mathcal{F}$ is used as a critique that would provide a training signal to $\mathbf{V}$ to generate accurate foreground masks that would place $\mathcal{P}_{\mathrm{mask}}$ and $\mathcal{S}$ in close proximity in the latent space; otherwise a high loss is imposed. Given a distance function $\delta(\cdot,\cdot)$ the SBIR loss becomes:
$\mathcal{L}_{\mathrm{SBIR}}=\delta(\mathcal{F}(\mathcal{P}_{\mathrm{mask}}),\mathcal{F}(\mathcal{S}))\vspace{-0.1cm}$ (4)
Notably, for category-level and fine-grained segmentation, we use pre-trained category-level and fine-grained SBIR models [71] respectively.
Unpaired Sketch-Photo Loss.  Apart from paired sketch-photo data, we focus on unpaired photos to ensure that nothing is segmented (correlation map contains only 00s) from them, as they do not contain the sketched query concept. We implement this via unpaired sketch-photo loss, where, across each batch, for every query-sketch $\mathcal{S}$, we sample a random non-paired image $\mathcal{N}$ (from different instances for fine-grained, and different classes for category-level). We compute and reshape the patch-wise features of $\mathcal{N}$ as $\mathbf{n}_{s}$=$\mathbf{V}_{p}(\mathcal{N})\in\mathbb{R}^{h^{\prime}\times w^{\prime}\times d}$ and calculate cosine similarity between $\mathbf{n}_{s}$ and $\mathbf{s}$, followed by $\mathtt{Sigmoid}$ normalisation to generate the correlation map $\mathcal{C}_{n}$. The model is constrained to predict an all-zero $\mathcal{C}_{n}$ for random non-paired candidate images via a binary cross-entropy loss as:
$\mathcal{L}_{\mathrm{unpaired}}=-\frac{1}{h^{\prime}w^{\prime}}\sum_{i=1}^{h^{\prime}}\sum_{j=1}^{w^{\prime}}\mathtt{log}(1-\mathcal{C}_{n}(i,j)))\vspace{-0.3cm}$ (5)
Mask Regularisation Loss.  Ideally, we want to segment only the foreground sketch query concept suppressing the background pixels. Nonetheless, minimising $\mathcal{L}_{\mathrm{SBIR}}$ might result in a trivial solution of classifying all pixels as foreground object [89]. To eliminate the chance of a trivial mask with all $1$s, we utilise a regularisation loss that computes the cross-entropy between the generated mask $\mathcal{M}\in\mathbb{R}^{h\times w\times 1}$ and an all-zero mask as:
$\mathcal{L}_{\mathrm{reg}}=-\frac{1}{hw}\sum_{i=1}^{h}\sum_{j=1}^{w}\mathtt{log}(1-\mathcal{M}(i,j)))\vspace{-0.15cm}$ (6)
Thus, our overall training objective (Fig. 3) becomes $\mathcal{L}_{\mathrm{total}}=\lambda_{1}\mathcal{L}_{\mathrm{InfoNCE}}+\lambda_{2}\mathcal{L}_{\mathrm{SBIR}}+\lambda_{3}\mathcal{L}_{\mathrm{unpaired}}+\lambda_{4}\mathcal{L}_{\mathrm{reg}}$.
4.5 Further Extension to Part-level Segmentation
Part-level segmentation aims to restrict the model to segment only the region corresponding to the part-level sketch query (e.g., “head” of a camel). Unfortunately, common sketch-photo datasets [73] neither hold part-level sketches, nor part-level masks. We therefore aim to model this part-level behaviour via a simple sketch augmentation trick where we synthetically split the input sketch into multiple non-overlapping portions. Specifically, we first calculate the centroid of the query sketch from its vector coordinates. Now, at every pass, we draw a straight line from the centroid with a random slope $m_{1}$. Based on the number of non-overlapping portions $n$, we calculate the slope values ($m_{2},m_{3},\cdots,m_{n}$) for the next lines as $m_{n}$=$m_{n-1}+\frac{360}{n}$. These $n$ straight lines segment the sketch into $n$ orthogonal non-overlapping parts (Fig. 4). With $n$=$2$ (decided empirically), we segment the query sketch $\mathcal{S}$ into two non-overlapping parts $\mathcal{S}_{A}$ and $\mathcal{S}_{B}$. Using sketch-guided mask generation procedure (Sec. 4.3), we generate the corresponding masks $\mathcal{M}_{A}$ and $\mathcal{M}_{B}$. Ideally, $\mathcal{M}_{A}$ and $\mathcal{M}_{B}$ should only segment their corresponding regions from the candidate images in a non-overlapping manner, which therefore incurs a penalisation if $\mathcal{M}_{A}$ and $\mathcal{M}_{B}$ overlap. To do so we combine $\mathcal{M}_{A}$ and $\mathcal{M}_{B}$ via bit-wise XOR to get the final mask as $\mathcal{M}$=$\mathcal{M}_{A}\oplus\mathcal{M}_{B}$, where the common foreground regions should be nullified. Consequently, the same regions will be suppressed in the masked candidate image $\mathcal{P}_{\mathrm{mask}}$=$\mathcal{P}\times{\mathcal{M}}$ as well. Accordingly, $\mathcal{L}_{\mathrm{SBIR}}$ would impose a high loss, leading to better part-specific segmentation.
Figure 4: Example of our sketch-partitioning augmentation. $\mathcal{S}$ is divided into $\mathcal{S}_{A}$ and $\mathcal{S}_{B}$ based on the straight lines from centroid (blue dot). Common foreground region of $\mathcal{M}$ is bordered in red.
4.6 Discussion
For category-level segmentation, we use a pre-trained category-level SBIR model (Sec. 3) with a pre-trained CLIP [68] ViT-L/14 backbone due to its category-level generalisation potential [68]. Whereas, for fine-grained segmentation, we use a pre-trained FG-SBIR model (Sec. 3) with a pre-trained DINOv2 [60] ViT-L/14 backbone. For extension to part-level, we use the same fine-grained segmentation model with the sketch-partitioning augmentation (Sec. 4.5) applied to $50\%$ of sketch samples in each batch.
During testing, $\mathbf{V}_{g}$ and $\mathbf{V}_{p}$ extract global sketch feature $\mathbf{s}$ and patch-wise candidate image features $\mathbf{p}_{s}$, using which we calculate the sketch-guided correlation map $\mathcal{C}$ (Eq. 2). Upscaling $\mathcal{C}$, followed by hard thresholding it with a fixed value of $0.5$ yields the final segmentation mask $\mathcal{M}$, which then undergoes standard CRF [42]-based post-processing.
Methods Category-level Fine-grained
mIoU (S) mIoU (U) hIoU pAcc. mIoU (S) mIoU (U) hIoU pAcc.
Baselines S-B1 60.5 51.8 54.7 61.5 50.7 41.8 42.5 57.8
S-B2 67.7 55.2 60.4 69.7 51.3 46.1 47.6 61.6
S-B3 56.2 47.6 51.2 58.8 49.1 40.9 42.7 50.5
T-B1 66.3 56.7 57.9 70.2 50.2 42.7 43.2 61.7
Adapted SOTAs S-PerSAM 67.7 47.6 53.2 69.9 54.2 40.6 47.1 60.6
S-ZS-Seg 68.2 56.9 54.5 71.1 56.4 48.8 49.6 67.4
S-Graph-Cut 56.9 46.2 50.4 56.6 44.1 36.7 39.8 47.2
S-Sketch-a-Segmenter 68.9 57.3 56.1 72.4 57.5 49.1 51.4 68.2
Ours 70.1 61.3 59.7 76.8 66.9 56.2 52.8 74.7
Table 1: Quantitative comparison on the Sketchy-Extended [48] and Sketchy [73] for category-level and fine-grained segmentation.
Figure 5: Qualitative comparison on Sketchy-Extended [48] for category-level segmentation on seen (left) and unseen (right) classes.
Figure 6: Qualitative comparison on Sketchy [73] for fine-grained segmentation on seen (top) and unseen (bottom) classes.
5 Experiments
Dataset.   As our method does not require ground truth masks, we can utilise any standard sketch-photo dataset for training. Here, we use Sketchy [73] and Sketchy-extended [48, 73] to train our category-level and fine-grained models respectively. While Sketchy [73] holds $12,500$ photos from $125$ classes, each having at least $5$ fine-grained paired sketches, its extended version [48] carries additional $60,652$ photos from ImageNet [23].
Evaluation.   We evaluate for both category-level (Sketchy-extended) and fine-grained (Sketchy) setups individually on two paradigms – (i) for unseen set evaluation, we train our method on $104$ Sketchy classes and test on the $21$ unseen classes; (ii) for seen set evaluation, we randomly divide the $104$ seen-classes into $80$:$20$ train:test splits across samples. To evaluate segmentation accuracy, we manually annotate each test image from the Sketchy dataset with segmentation label masks to form the $\mathtt{SketchySegment}$ dataset. For further research, we open-sourced the $\mathtt{SketchySegment}$ dataset (cf§ Supplementary). Notably, our training pipeline does not involve any ground truth masks. Following standard literature [96, 85, 57], we use mean intersection over union (mIoU), and per-pixel classification accuracy (pAcc.) as our primary evaluation metrics for both seen and unseen sets. Additionally, we also report the harmonic mean IoU (hIoU) among the unseen and seen sets.
Implementation Details.  We use a pre-trained ViT-L/14 CLIP [68] and ViT-L/14 DINOv2 [60] backbone for category-level and fine-grained segmentation models respectively. We fine-tune the $\mathtt{LayerNorm}$ layers of the CLIP visual encoder with a learning rate of $10^{-5}$, keeping the rest of the parameters frozen. The DINOv2 encoder is fine-tuned with a low learning rate of $10^{-6}$. We train the model for $200$ epochs, in an NVIDIA RTX 2080-Ti GPU using AdamW [50] optimiser and a batch size of $16$. All $\lambda$ values are set to $1$, empirically. We use pre-trained baseline SBIR and FG-SBIR models from [71].
[…]
5.1 Performance Analysis
Category-level Segmentation.Tab. 1 shows the following observations – (i) although S-B2 outperforms S-B1 (due to sub-optimal embedding-space alignment) and S-B3 (owing to lack of mask regularisation), all these baselines fare poorly when compared to ours, particularly for S-B3 ($19.82\%$ mIoU (S) drop vs. ours), justifying the necessity of mask regularisation via $\mathcal{L}_{\mathrm{reg}}$. (ii) S-ZS-Seg with its explicit zero-shot modelling [85] depicts comparable performance on both seen and unseen classes. However, S-PerSAM despite leveraging the large-scale pre-training of SAM [39], turns out to be sub-optimal for the unseen classes. (iii) owing to the inferior pseudo-mask generation, S-Graph-Cut fails to beat even the stronger baselines (i.e., S-B2, T-B1). Our method outperforms these competitors with an average $8.63~{}(14.49)\%$ mIoU gain on the seen (unseen) classes, even without any ground truth masks during training. Qualitative comparisons are shown in Fig. 5.
Fine-grained Segmentation.  Fine-grained setup in Tab. 1 shows: (i) although T-B1 competes ours in category-level setup, it scores a lower mIoU in the more challenging fine-grained setup. We posit that, while simple handcrafted prompts (i.e., $\mathtt{a~{}photo~{}of~{}a~{}[CLASS]}$) of T-B1 sufficed the category-level task, it failed to capture the fine-grained shape/structure. Although adding more descriptions might improve performance for T-B1, the research question we raise is: how fine-grained a textual description can be? Fine-grained captions generated by SOTA captioners [43, 44] often suffer from noise and inaccuracy [43, 44]. Additionally, crafting detailed textual prompts is more laborious compared to freehand sketching [8, 10, 19]. (ii) despite being a stronger category-level baseline, S-B2 performs sub-optimally ($23.31\%$ mIoU (S) drop vs. ours) here, verifying the need for $\mathcal{L}_{\mathrm{unpaired}}$ in fine-grained setting. (iii) our method outperforms all sketch-adapted SOTAs with an average hIoU margin of $13.84\%$ without the complicated positive-negative prior calculation of S-PerSAM, time-consuming two-stage approach of S-ZS-Seg, or the allegedly unstable pseudo-mask generation of S-Graph-Cut. Qualitative comparisons (Fig. 6) show the fine-grained generalisation potential of our method. (iv) S-Sketch-a-Segmenter despite being fully-supervised, achieves lower hIoU and pAcc. in the Sketchy dataset.
[…]
Figure 7: Part-level segmentation: T-B1 vs. Ours on Sketchy [73].
Segmenting Composed Concepts.  We also test our framework for composed concept segmentation. For instance, drawing a “dog” sitting beside a “frisbee” should segment both the dog and the frisbee. Qualitative comparison (on FS-COCO [58]) in Fig. 8 shows our method to perform reasonably better in segmenting composed concepts than T-B1.
Figure 8: Composed segmentation: T-B1 vs. Ours.
5.2 Ablation on Design
[i] Backbone Variants.  Performance of our sketch-based image segmenter depends crucially on the choice of feature extractor backbone. Accordingly, we explore with multiple publicly available pre-trained DINOv2 [60] and CLIP [68] backbones. Quantitative results in Tab. 2, show DINOv2 ViT-L/14 and CLIP ViT-L/14 to achieve the highest mIoU for fine-grained and category-level segmentation respectively. Although CLIP [68] and DINOv2 [60] showcase comparable category-level performance, we posit that, due to its self-supervised pre-training, DINOv2 [60] outperforms CLIP [68] in the fine-grained setup.
[ii] Does Fine-tuning entire CLIP help?  Existing literature on sketch-based vision tasks [74, 72] point towards fine-tuning the entire CLIP vision encoder to adapt it to their respective tasks [74, 72]. Doing so for training our category-level segmentation model however, significantly drops the final seen/unseen class mIoU (CLIP Fine-tuning result in Tab. 2). We hypothesise that fine-tuning all parameters of the vision encoder distorts the pre-trained knowledge of CLIP, thus sacrificing its generalisation potential.
[iii] Effect of $\mathcal{L}_{\mathrm{InfoNCE}}$.  We posit that the InfoNCE loss aids in proper alignment of sketch-photo features to generate accurate patch-wise correlation maps. Hence, a drop of $12.69~{}(17.48)\%$ seen-class mIoU in w/o $\mathcal{L}_{\mathrm{InfoNCE}}$ (Tab. 2), for category-level (fine-grained) setup, verifies its importance in learning patch-wise sketch-photo region alignment.
[iv] Why Sketch-Partitioning Augmentation?   Quantitative assessment of sketch-partitioning for part-level segmentation is infeasible due to the lack of annotation masks. However w/o Sketch-partitioning results in Tab. 2 shows our part-level augmentation to be particularly helpful for our fine-grained segmentation model, as its absence further reduces the seen (unseen) class mIoU by $8.22~{}(10.32)\%$.
Methods Category-level Fine-grained
mIoU (S) mIoU (U) mIoU (S) mIoU (U)
DINOv2 ViT-S/14 58.7 50.9 60.4 50.3
ViT-B/14 65.1 57.2 63.6 51.7
ViT-L/14 68.5 59.8 66.9 56.2
CLIP ViT-B/16 59.7 51.2 54.6 43.3
ViT-B/32 66.4 59.9 56.4 44.8
ViT-L/14 70.1 61.3 60.1 49.7
w/o $\mathcal{L}_{\text{InfoNCE}}$ 61.2 52.8 55.2 48.1
w/o Sketch-partitioning 69.3 60.1 61.4 50.4
CLIP Fine-tuning 60.6 32.7
Ours-full 70.1 61.3 66.9 56.2
Table 2: Ablation on design.
5.3 Limitations and Failure Cases
Although our method is quite robust against deformed sketches and hard-to-segment candidate images (Fig. 5-8), it sometimes struggles to yield accurate segmentation masks in case of challenging candidate images (Fig. 9) with reflection, cluttered background, occlusion, or camouflage.
Figure 9: Failure cases. (Best view when zoomed in.)
6 Conclusion
Here, we explore the potential of freehand abstract sketches for subjective image segmentation. Leveraging large-scale pre-trained visual encoders and SBIR models, we devise a mask-free image segmenter that allows for segmentation at multiple granularity levels (i.e., category-level and fine-grained). Furthermore, using the smart augmentation trick of sketch-partitioning, we extend our method to part-level segmentation. Extensive qualitative and quantitative results depict the efficacy of our method. We hope that this study paves the path for more sketch-based vision tasks in future.

### v2

@@
-Freestyle Sketch-in-the-Loop Image Segmentation
+SketchYourSeg: Mask-Free Subjective Image Segmentation via Freehand Sketches
 Abstract
-In this paper, we expand the domain of sketch research into the field of image segmentation, aiming to establish freehand sketches as a query modality for subjective image segmentation.
-Our innovative approach introduces a “sketch-in-the-loop” image segmentation framework, enabling the segmentation of visual concepts partially, completely, or in groupings – a truly “freestyle” approach – without the need for a purpose-made dataset (i.e., mask-free).
-This framework capitalises on the synergy between sketch-based image retrieval (SBIR) models and large-scale pre-trained models (CLIP or DINOv2).
-The former provides an effective training signal, while fine-tuned versions of the latter execute the subjective segmentation.
-Additionally, our purpose-made augmentation strategy enhances the versatility of our sketch-guided mask generation, allowing segmentation at multiple granularity levels.
-Extensive evaluations across diverse benchmark datasets underscore the superior performance of our method in comparison to existing approaches across various evaluation scenarios.
+We introduce SketchYourSeg, a novel framework that establishes freehand sketches as a powerful query modality for subjective image segmentation across entire galleries through a single exemplar sketch.
+Unlike text prompts that struggle with spatial specificity or interactive methods confined to single-image operations, sketches naturally combine semantic intent with structural precision.
+This unique dual encoding enables precise visual disambiguation for segmentation tasks where text descriptions would be cumbersome or ambiguous – such as distinguishing between visually similar instances, specifying exact part boundaries, or indicating spatial relationships in composed concepts.
+Our approach addresses three fundamental challenges: (i) eliminating the need for pixel-perfect annotation masks during training with a mask-free framework; (ii) creating a synergistic relationship between sketch-based image retrieval (SBIR) models and foundation models (CLIP/DINOv2) where the former provides training signals while the latter generates masks; and (iii) enabling multi-granular segmentation capabilities through purpose-made sketch augmentation strategies.
+Our extensive evaluations demonstrate superior performance over existing approaches across diverse benchmarks, establishing a new paradigm for user-guided image segmentation that balances precision with efficiency.
 1 Introduction
-This paper delves into the realm of image segmentation, but with a unique twist – the incorporation of freehand sketches.
-Our research vision is to explore the impact of fine-grained subjectivity inherently captured by a sketch [30, 19], on image segmentation.
-At its core, our aim is to devise a subjective image segmentation framework that empowers users to freely segment any object, be it a partial segment, a complete entity, or a group of interconnected elements – essentially, enabling image segmentation in a “freestyle” manner.
-Image segmentation has witnessed significant evolution [14, 39] over time, and we are committed to keeping pace with the current capabilities of image segmentation systems – we also want to operate in an open-vocabulary [57] manner and harness the potential of the latest foundation models.
-We too fully acknowledge prior segmentation methods that inject subjectivity through textual inputs [85].
-The central argument of this paper however revolves around the premise that sketches offer a more precise form of subjectivity – a point convincingly illustrated in LABEL:fig:teaser (bottom).
-To elucidate why sketches serve as a superior modality for capturing subjectivity, we direct readers to the extensive sketch literature [10, 19, 86, 25, 74].
-In summary, the insights drawn from this large body of work are two-fold: (i) sketches excel in their ability to convey intricate visual details that may be cumbersome or ambiguous when expressed through text [74].
-(ii) sketches exhibit a smaller semantic gap to photos, as both share the same visual modality, albeit in an abstract form [30].
-This inherent relationship is notably featured in the extensively studied domain of SBIR [20] and its fine-grained variations [7] – a challenge that we also employ as a guiding critic in our training.
-The specifics of our vision present significant challenges.
-First and foremost, we need to operate without any sketch-mask pairing data.
-The collection of such a dataset is simply unfeasible, especially considering the current status quo within the sketch community still concerns gathering enough sketch-photo pairs [58, 73], let alone pixel-perfect masks.
-This challenge is further exacerbated by our open-vocabulary setting, which essentially renders prior arts inapplicable [31].
-Secondly, akin to most prior work in the sketch literature [86, 72, 6, 41], we face a substantial gap between the sketch and photo domains.
-We commence by fixating on the ultimate goal: given a sketch, compute corresponding segmentation masks for all applicable gallery photos.
-Surprisingly, it turns out that all we need to compute this correspondence is the strategic utilisation of (i) a frozen Fine-Grained Sketch-Based Image Retrieval (FG-SBIR) model and (ii) a pre-trained foundation model (e.g., CLIP [68], DINOv2 [60]).
-The former provides a solid foundation for fine-grained sketch-photo correspondence and also serves as a training signal to fine-tune the latter.
-Meanwhile, the fine-tuned version of the latter predicts the actual mask for the photo using its corresponding sketch, thereby eliminating the need for manually annotated masks (i.e., mask-free).
-It follows that these two components are intricately linked by this very hypothesis: a sketch and its matching background-masked image pair (and foreground highlighted) will sit closely within the fixed FG-SBIR joint embedding [72].
-This, we hope, is intuitive – the more highlighted (using the computed mask) the foreground object is, the easier it is for its matching sketch to locate.
-It is through this iterative training process that the segmentation model continually hones its performance.
-In generating sketch-guided masks, we use pre-trained DINOv2 [60] for the fine-grained setup (CLIP [68] for category-level).
-It extracts a spatial feature map from a photo and a global feature vector from a sketch.
-A cosine similarity computation, followed by bilinear upscaling and differentiable thresholding, yields a pseudo mask used for generating the background-masked image corresponding to the given sketch.
-To prevent all-white mask predictions and overfitting, a mask-regularisation loss is introduced.
-Additionally, for part-level segmentation, we employ sketch-partitioning, randomly fragmenting binary sketches into orthogonal splits, enabling model-prediction from a specific split, for more precise part-level segmentation.
-In summary, our contributions are (i) we address the challenge of subjective segmentation by introducing a novel approach that enables mask-free image segmentation through the interpretive power of sketches.
-(ii) we support multi-granularity segmentation, providing users with the flexibility to segment visual concepts at various levels, encompassing entire entities, specific parts, or interconnected groups.
-(iii) we establish an effective synergy between sketch-based image retrieval models and large-scale pre-trained models.
-(iv) we propose a purpose-made augmentation strategy that enhances the framework’s versatility, enabling segmentation at multiple granularity levels.
+This paper introduces a novel approach to image segmentation that leverages freehand sketches for subjective segmentation across entire galleries with a single exemplar sketch.
+Our method, SketchYourSeg, enables users to express personalised segmentation intent through simple sketches rather than complex text descriptions or repetitive interactive inputs, whether targeting whole objects, specific parts, or multiple related elements.
+Image segmentation has evolved substantially [36], with current systems increasingly leveraging foundation models in open-vocabulary settings [55].
+However, a critical question remains: how can users effectively communicate their subjective segmentation intent while efficiently operating across multiple images?
+While text prompts [85] and interactive methods [16, 26] have gained popularity, they present fundamental limitations for precisely specifying segmentation targets, especially across entire image sets.
+To address this challenge, we investigate freehand sketches as a query modality for subjective segmentation.
+A key question arises: why use sketches when alternatives like text prompts or interactive scribbles exist?
+Research evidence demonstrates that sketches occupy a unique middle ground in the specificity-effort tradeoff curve.
+Text prompts, while easy to produce and capable of gallery-wide operations, struggle with spatial and structural specificity for fine-grained tasks [39, 56, 17, 75, 9].
+For instance, precisely describing “the zebra with distinctive stripes that curve inward at the shoulder” through text alone presents significant challenges that sketches can naturally overcome.
+Interactive methods using clicks or scribbles, while offering precise spatial control, operate only on one image at a time and typically require multiple interactions per image [80].
+This fundamental limitation makes them impractical for gallery-wide operations where users need to segment the same concept across multiple images simultaneously.
+Sketches, by contrast, naturally combine semantic intent with structural specificity in a single interaction [40].
+This dual encoding enables our approach to segment visual concepts across entire image galleries through a single exemplar sketch – a capability neither text nor interactive methods can match.
+The sketch serves as both a semantic concept descriptor (similar to text) and a structural template (similar to scribbles), making it uniquely suited for fine-grained segmentation tasks where spatial relationships and distinctive visual features play a crucial role.
+Our approach confronts two significant technical challenges.
+First, we must operate without sketch-mask pairing data, as collecting such a dataset is infeasible given the sketch community’s ongoing struggle to gather even basic sketch-photo pairs [56, 71].
+Second, we must bridge the substantial domain gap between abstract sketches and photos [86, 70] at the pixel level.
+We address these challenges through a novel mask-free approach that strategically utilizes: (1) a frozen Fine-Grained Sketch-Based Image Retrieval (FG-SBIR) model, and (2) a pre-trained foundation model (CLIP [66] or DINOv2 [58]).
+These components are linked by our central hypothesis: a sketch and its matching background-masked image will occupy proximate positions within a fixed FG-SBIR embedding space [70].
+This creates a self-supervised loop – the better our segmentation mask highlights the relevant content, the closer the masked image and sketch will be in the embedding space, providing a natural training signal without manual annotations.
+For mask generation, we extract spatial feature maps from photos and global feature vectors from sketches.
+Using cosine similarity, bilinear upscaling, and differentiable thresholding, we generate pseudo masks that highlight regions corresponding to the sketch query.
+To prevent trivial solutions and enable multi-granular segmentation, we introduce mask regularisation techniques and a novel sketch-partitioning strategy that enables part-level segmentation without additional training data.
+Our primary contributions are: (i) A mask-free training paradigm that eliminates the need for pixel-perfect annotations while still achieving high-quality segmentation results.
+(ii) A synergistic framework that leverages sketch-based image retrieval models as critics to guide foundation models in generating accurate segmentation masks.
+(iii) Multi-granular segmentation capabilities that leverage sketches’ unique structural encoding to precisely specify visual intent that text cannot effectively communicate – from shape-specific distinctions between similar instances to exact part boundaries and spatial relationships between objects – all within a single framework.
+Empirical validation through extensive benchmark evaluations, demonstrating superior performance over existing approaches for both seen and unseen categories.
 2 Related Works
@@
 To our best knowledge, Sketch-a-Segmenter [28] is the only existing work dealing with sketch-guided segmentation.
-This method [31] however requires fully-annotated pixel-level masks for training [31], unlike ours.
-We thus here aim to forward sketch research in another interesting direction by utilising its potential as a weak label for segmentation.
 Segmentation in Vision.
@@
 might suffice for category-level vision tasks, they lack in fine-grained cues unlike sketches [9, 18].
-Inspired by sketch’s success in multiple fine-grained vision tasks (e.g., retrieval [72], detection [19], generation [41], etc.), we aim to establish sketch as a worthy query-modality for mask-free weakly-supervised image segmentation, both at category-level and fine-grained setup.
-Furthermore, we also extend our method to part-level segmentation, using the same dataset.
+Inspired by sketch’s success in multiple fine-grained vision tasks (e.g., retrieval [70], detection [18], generation [38], etc.), we aim to establish sketch as a worthy query-modality for mask-free weakly-supervised image segmentation, at category-level, fine-grained, and part-level segmentation setup.
 4.1 Problem Definition: Segmentation Granularity
@@
 4.2 Baseline Weakly Supervised Segmentation
-In our setup, we consider the available sketches ($\mathcal{S}$) and photos ($\mathcal{P}$) from $\mathcal{D}_{\mathrm{cat}}$= $\{\mathcal{S}_{i},\mathcal{P}_{i}\}_{i=1}^{M}$ and $\mathcal{D}_{\mathrm{fine}}$=$\{\mathcal{S}_{j},\mathcal{P}_{j}\}_{j=1}^{N}$ with category-level [48] and fine-grained [73] association between them respectively (cf.
-Sec.
-3).
-We design our baseline sketch-based image segmenter taking inspiration from text-based weakly-supervised segmentation framework [57] that employs patch-wise InfoNCE loss between CLIP’s global textual feature and patch-wise visual features [57] for better text-photo region-wise alignment (instead of using global feature for both as in CLIP training [68]) required for dense pixel-level segmentation task [57].
+We consider the available sketches ($\mathcal{S}$) and photos ($\mathcal{P}$) from $\mathcal{D}_{\mathrm{cat}}$=$\{\mathcal{S}_{i},\mathcal{P}_{i}\}_{i=1}^{M}$ and $\mathcal{D}_{\mathrm{fine}}$=$\{\mathcal{S}_{j},\mathcal{P}_{j}\}_{j=1}^{N}$ with category-level and fine-grained association between them respectively.
+We design the baseline sketch-based image segmenter taking inspiration from text-based weakly-supervised segmentation method [55] that uses patch-wise InfoNCE loss between CLIP’s global textual feature and patch-wise visual features [55] for better text-photo region-wise alignment (instead of using global feature for both as in CLIP training) required for dense pixel-level segmentation task.
 We however exclude the textual encoder and employ CLIP’s visual encoder to extract patch-wise photo features and global query sketch feature to apply the same loss.
@@
 We get the sketch-guided correlation map $\mathcal{C}$ by computing cosine similarity between $\mathbf{s}$ and $\mathbf{p}_{s}$ using the sketch as a probe-vector as:
-$\small\mathcal{C}(u,v)=\frac{\langle\mathbf{s},\mathbf{p}_{s}(u,v)\rangle}{||\mathbf{s}||~{}||\mathbf{p}_{s}(u,v)||}~{}~{}~{}~{}u=[1,..,h^{\prime}],v=[1,..,w^{\prime}]\vspace{-0.1cm}$ (2)
+$\mathcal{C}(u,v)=\frac{\langle\mathbf{s},\mathbf{p}_{s}(u,v)\rangle}{||\mathbf{s}||~{}||\mathbf{p}_{s}(u,v)||}~{}~{}~{}~{}u=[1,..,h^{\prime}],v=[1,..,w^{\prime}]\vspace{-0.1cm}$ (2)
 Now the research question remains as to how can we leverage this sketch-guided correlation map $\mathcal{C}$ in the learning process while backpropagating the training signal via pixel space.
@@
 4).
+Figure 4: Example of our sketch-partitioning augmentation.
+$\mathcal{S}$ is divided into $\mathcal{S}_{A}$ and $\mathcal{S}_{B}$ based on the straight lines from centroid (blue dot).
+Common foreground region of $\mathcal{M}$ is bordered in red.
 With $n$=$2$ (decided empirically), we segment the query sketch $\mathcal{S}$ into two non-overlapping parts $\mathcal{S}_{A}$ and $\mathcal{S}_{B}$.
@@
 Accordingly, $\mathcal{L}_{\mathrm{SBIR}}$ would impose a high loss, leading to better part-specific segmentation.
-Figure 4: Example of our sketch-partitioning augmentation.
-$\mathcal{S}$ is divided into $\mathcal{S}_{A}$ and $\mathcal{S}_{B}$ based on the straight lines from centroid (blue dot).
-Common foreground region of $\mathcal{M}$ is bordered in red.
 4.6 Discussion
@@
 S-Graph-Cut 56.9 46.2 50.4 56.6 44.1 36.7 39.8 47.2
-S-Sketch-a-Segmenter 68.9 57.3 56.1 72.4 57.5 49.1 51.4 68.2
 Ours 70.1 61.3 59.7 76.8 66.9 56.2 52.8 74.7
@@
 The DINOv2 encoder is fine-tuned with a low learning rate of $10^{-6}$.
-We train the model for $200$ epochs, in an NVIDIA RTX 2080-Ti GPU using AdamW [50] optimiser and a batch size of $16$.
+We train the model for $200$ epochs using AdamW [49] optimiser and a batch size of $16$.
 All $\lambda$ values are set to $1$, empirically.
@@
 2).
-We hypothesise that fine-tuning all parameters of the vision encoder distorts the pre-trained knowledge of CLIP, thus sacrificing its generalisation potential.
+We posit that fine-tuning all parameters of the vision encoder distorts the pre-trained knowledge of CLIP, thus sacrificing its generalisation potential.
 [iii] Effect of $\mathcal{L}_{\mathrm{InfoNCE}}$.
 We posit that the InfoNCE loss aids in proper alignment of sketch-photo features to generate accurate patch-wise correlation maps.
-Hence, a drop of $12.69~{}(17.48)\%$ seen-class mIoU in w/o $\mathcal{L}_{\mathrm{InfoNCE}}$ (Tab.
-2), for category-level (fine-grained) setup, verifies its importance in learning patch-wise sketch-photo region alignment.
+Hence, a drop of $12.69~{}(17.48)\%$ seen-class mIoU in w/o $\mathcal{L}_{\mathrm{InfoNCE}}$, for category-level (fine-grained) setup, verifies its importance in learning patch-wise sketch-photo region alignment.
 [iv] Why Sketch-Partitioning Augmentation?
@@
 Extensive qualitative and quantitative results depict the efficacy of our method.
-We hope that this study paves the path for more sketch-based vision tasks in future.